# Optimizing Language Models for Inference Time Objectives using Reinforcement Learning

### v1

@@ 3 Optimizing Language Models for Inference Time Objectives using RL
defined through a function $f:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}...\rightarrow\mathbb{R}$. Here $f$ is an aggregation function that can process an arbitrary number of generations. It is clear that both pass@$k$ and majority voting objectives are special cases.
@@ 3.2 Trading-off further variance reduction with bias
The objective differs from the initial $k$-sample objective in Eqn (1) - it measures the leave-one-out $k-1$-sample objective $f\left(\mathbf{y}_{-i}\right)$ averaged over all $k$ samples. The bias is hence
@@ 4 Related work
The leave-one-out control variate has been studied in Mnih & Rezende (2016) for variational inference, and in Kool et al. (2019); Ahmadian et al. (2024) for improving the mean objective in a REINFORCE algorithm. As an alternative, Mnih & Rezende (2016) analyzed the control variate $v_{i}\coloneqq\mathbb{E}_{\mathbf{y}_{-i}}\left[f(\mathbf{y})|y_{i}\right]$ which accounts for the credit assigned to other $k-1$ samples in an expected sense. Such an objective can be especially useful for the $k$-sample objectives in our case since the advantage $\tilde{A}_{i}=f(\mathbf{y})-v_{i}$ is by construction zero-mean and arguably has lower variance. However, estimating such quantities introduces additional overhead in practice.
@@ 5 Experiments on mathematical reasoning
Throughout, we focus on the mathematical reasoning dataset MATH (Hendrycks et al., 2021) where the prompt $x$ consists in asking the model a mathematical question with a short-form ground truth answer $a^{*}$ available. Given the model generation $y=(c,a)$ which typically consists of a step-by-step chain-of-thought reasoning $c$ and a proposed answer $a$, the reward is computed as a match between $a$ and $a^{\ast}$. We adopt Sympy (Meurer et al., 2017) to automatically match the answers and assign a reward of $r=1$ if there is a match and $r=-1$ otherwise. As such, the objective $\mathbb{E}[r]$ also measures the average accuracy of the policy. We train on models on the MATH training set with various objective alternatives introduced above. During experiments, we provide the model with a system prompt that asks for a step-by-step solution, followed by a final answer.
@@ 6 Experiments on code generation tasks
For code generation, we conduct our experiments on CodeContests (Li et al., 2022), a competitive programming benchmark containing 13k problems as training set and evaluate on the valid and test set. Similar to previous work (Xu et al., 2024; Gehring et al., 2024), we use a $r=+1/-1$ reward indicating whether the Python code solutions pass all the given tests that come with the dataset. Due to the challenging nature of CodeContests, we use Llama 3.1 70B Instruct model and optimize for pass@8 metrics on the CodeContests training set. To measure generalization, we also report the performance on another competitive programming benchmark, TACO (Li et al., 2023) on the easy and hard split, for which we have decontaminated the CodeContests training set against TACO eval set.
@@
We also investigate how the $k$-sample gradient variants can be combined with much more sophisticated algorithmic stack to illustrate its generic practical utility. Note that the pass@$k$ objective can be seamlessly integrated into other policy gradient variants, such as PPO (Schulman et al., 2017), which adds clipping and an additional value function that serves as the baseline. Figure 6 shows the performance of using PPO. We can observe that PPO with pass@1 objective enters in high-KL regime at the end of the training, due to the fact that its update generally leverages much more dense signals. The sparse nature of pass@$k$ objective keeps the policy under low-KL regime and maintains higher pass@8 performance. For the purpose of training language models with fixation on the reference policy (Ziegler et al., 2019; Ouyang et al., 2022), pass@$k$ objectives achieve a balance for the performance metrics.
@@ 6.1 Generalization to different $k$ for evaluation
We show in Table 1 and Table 2 (in the Appendix) that Llama 3.1 70B Instruct model trained to optimize pass@$k$ ($k$ = 8) can generalize to pass rate with different $k$ up to $k=100$, with over $+10\%$ solve rate improvement over the baseline. on validation and test set.
@@ 7 Discussion and limitations
With a focus on the pass@$k$ and majority voting inference objectives, we have demonstrated a clear trade-off: for specific inference time objectives, it is typically helpful to optimize for those objectives with specialized losses at training time. Mechanistically, such objectives might strike for an empirical trade-off important to the downstream inference algorithm, such as the mean performance and sample diversity. Naturally, the improvement depends on the relative capability of the base model and the problems of interest - in cases inference time objectives and mean performance are largely aligned, the mean policy gradient algorithm remains a strong baseline, an useful observation for practitioners.
@@ Appendix A Hyper-parameters and experimental details
[…]
@@ A.1 Ablation on number of samples $k$
A.2 Training on HARP
@@ A.3 Evaluation
A.4 Experimental details and additional result on CodeContests and TACO
Experimental details.
[…]
Integrating with PPO.
$J^{\pi}(\theta)$ $=\mathbb{\hat{E}}_{y\sim\pi_{\theta_{\text{old}}}}\left[\min\left(\frac{\pi_{\theta}(y|x)}{\pi_{\theta_{\text{old}}}(y|x)}\hat{A},\operatorname{clip}\left(\frac{\pi_{\theta}(y|x)}{\pi_{\theta_{\text{old}}}(y|x)},1-\epsilon,1+\epsilon\right)\hat{A}\right)\right]$
$\hat{A}_{i}$ $=R_{i}-V_{\psi}(x),$
Mean(pass@1) training objective: $\quad R_{i}=r_{i}$
with $r_{i}$ being the +1/1+1/-1+ 1 / - 1 binary reward of whether the code $y_{i}$ is correct. We train the value model $V_{\psi}$ to minimize the following clipped value loss:
We set $\epsilon=\alpha=0.2$ in our experiments.
Llama 3.1 8B Instruct Performance.
@@ B.2 Interpolation between $k$-sample objectives and mean objective
where $S_{p}$ is the set of all subsets of $p$ indices from $\{1,\dots,k\}$, with $\left({k\atop p}\right)$ being its cardinality. Notice that for $p=0$ we recover the initial problem of the pass@$k$ objective. For $p=1$ we get the de-meaned objective introduced in the previous section. For $p=k-1$ we get the problem of maximizing the expected rewards. For $p=2$ we get the objective

### v2

@@ 3 Optimizing Language Models for Inference Time Objectives using RL
defined through a function $f:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}$$\rightarrow\mathbb{R}.$ Here $f$ is an aggregation function that can process an arbitrary number of generations. It is clear that both pass@$k$ and majority voting objectives are special cases.
@@ 3.2 Trading-off further variance reduction with bias
The objective differs from the initial $k$-sample objective in Eqn (1) - it measures the leave-one-out k1k-1-sample objective $f\left(\mathbf{y}_{-i}\right)$ averaged over all $k$ samples. The bias is hence
@@ 4 Related work
The leave-one-out control variate has been studied in Mnih & Rezende (2016) for variational inference, and in Kool et al. (2019); Ahmadian et al. (2024) for improving the mean objective in a REINFORCE algorithm. As an alternative, Mnih & Rezende (2016) analyzed the control variate $v_{i}\coloneqq\mathbb{E}_{\mathbf{y}_{-i}}\left[f(\mathbf{y})|y_{i}\right]$ which accounts for the credit assigned to other k1k-1 samples in an expected sense. Such an objective can be especially useful for the $k$-sample objectives in our case since the advantage $\tilde{A}_{i}=f(\mathbf{y})-v_{i}$ is by construction zero-mean and arguably has lower variance. However, estimating such quantities introduces additional overhead in practice.
@@ 5 Experiments on mathematical reasoning
Throughout, we focus on the mathematical reasoning dataset MATH (Hendrycks et al., 2021) where the prompt $x$ consists in asking the model a mathematical question with a short-form ground truth answer $a^{*}$ available. Given the model generation $y=(c,a)$ which typically consists of a step-by-step chain-of-thought reasoning $c$ and a proposed answer $a$, the reward is computed as a match between $a$ and $a^{\ast}$. We adopt Sympy (Meurer et al., 2017) to automatically match the answers and assign a reward of $r=1$ if there is a match and r=1r=-1 otherwise. As such, the objective $\mathbb{E}[r]$ also measures the average accuracy of the policy. We train on models on the MATH training set with various objective alternatives introduced above. During experiments, we provide the model with a system prompt that asks for a step-by-step solution, followed by a final answer.
@@ 6 Experiments on code generation tasks
For code generation, we conduct our experiments on CodeContests (Li et al., 2022), a competitive programming benchmark containing 13k problems as training set and evaluate on the valid and test set. Similar to previous work (Xu et al., 2024; Gehring et al., 2024), we use a r=+1/1r=+1/-1 reward indicating whether the Python code solutions pass all the given tests that come with the dataset. Due to the challenging nature of CodeContests, we use Llama 3.1 70B Instruct model and optimize for pass@8 metrics on the CodeContests training set. To measure generalization, we also report the performance on another competitive programming benchmark, TACO (Li et al., 2023) on the easy and hard split, for which we have decontaminated the CodeContests training set against TACO eval set.
@@
We also investigate how the $k$-sample gradient variants can be combined with much more sophisticated algorithmic stack to illustrate its generic practical utility. Note that the pass@$k$ objective can be seamlessly integrated into other policy gradient variants, such as PPO (Schulman et al., 2017), which adds clipping and an additional value function that serves as the baseline. Figure 6 shows the performance of using PPO. We can observe that PPO with pass@1 objective enters in high-KL regime at the end of the training, due to the fact that its update generally leverages much more dense signals. The sparse nature of pass@$k$ objective keeps the policy under low-KL regime and maintains higher pass@8 performance. For the purpose of training language models with fixation on the reference policy (Ziegler et al., 2019; Ouyang et al., 2022), pass@$k$ objectives achieve a balance for the performance metrics. We leave preliminary results of integrating with GRPO (Shao et al., 2024) in Appendix A.2.
@@ 6.1 Generalization to different $k$ for evaluation
We show in Table 1 and Table 2 (in the Appendix) that Llama 3.1 70B Instruct model trained to optimize pass@$k$ ($k$ = 8) can generalize to pass rate with different $k$ up to $k=100$, with over $10\%$ solve rate improvement over the baseline. on validation and test set.
@@ 7 Discussion and limitations
Impact Statement
This paper presents work whose goal is to advance fundamental algorithmic development. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Hyper-parameters and experimental details
We experimented with the Llama 3 model of size 3B, 8B and 70B (Dubey et al., 2024). All experiments are conducted with identical hyper-parameter settings: we always apply a batch size of $B=64$ prompts per update, and sample $k=4$ distinct generations per prompt by default. All training and evaluation sampling are conducted at a temperature of $\tau=1$ and with $\text{top-p}=1$.
@@ A.1 Ablation on number of samples $k$
A.2 Experimental details and additional result on CodeContests and TACO
Experimental details.
We experimented with Llama 3.1 8B and 70B Instruct model. For both model we use the same hyperparameters. We use a learning rate $2e^{-7}$, constant learning rate scheduling with $50$ warmup steps and weight decay of $0.1$. We sample $k=8$ generations per prompt. We update the model with a mini batch size 2 with sequence length $8192$ and train in total 8k gradient update steps. In training, sampling is conducted at a temperature of $\tau=1$ and with $\text{top-p}=1$. In evaluation, we sample at a temperature of $\tau=1$ and with $\text{top-p}=0.95$. Both pass@1 and pass@8 are estimated out of 20 samples. We evaluate the correctness of generated Python code using the official codebase of Li et al. (2022). Our asynchronous online training decouples the training and the inference, for which we include an importance sampling term to correct the slightly off-policy nature between current policy and the behavior policy. The original CodeContests training set contains 13328 problems. We follow the decontamination process presented by  Zheng et al. (2024) to decontaminate CodeContests training set against TACO evaluation set. We further remove problem instances with less than $5$ test cases. This results in total 12275 problems which we use to train our model.
Integrating with PPO.
[…]
$\hat{A}_{i}$ $=R_{i}-V_{\psi}(x),$
Mean(pass@1) training objective: $\quad R_{i}=r_{i}$
with $r_{i}$ being the +1/1+1/-1 binary reward of whether the code $y_{i}$ is correct. We train the value model $V_{\psi}$ to minimize the following clipped value loss:
We set $\epsilon=\alpha=0.2$ in our experiments.
Integrating with GRPO.
We can remove the baseline produced by the value model, $V_{\psi}(x)$, in the advantage estimation in the above PPO objective. The change is therefore only $\hat{A}_{i}=R_{i}$ and the rest remains unchanged. Interestingly, if we take the pass@$k$ training objective $R_{i}=\max_{j\in\{1,...,k\}}r_{j}-\max_{j\neq i}\ r_{j}$ and treat it as the “effective” reward, the biased pass@$k$ training objective coincides with the recently proposed Dr. GRPO (Liu et al., 2025) in which the baseline is the empirical mean. We can further divide it by the standard deviation to make it similar to the GRPO objective. We show the results of Llama 3.1 8B Instruct training on GRPO or Dr. GRPO objective in Table 3, with the normal pass@1 training (mean objective) as the baseline. The training has 8k gradient steps using the same set of hyperparameters. We report the number by averaging the performance of the last 5 checkpoints, i.e., checkpoints at 8k, 7.8k, …, 7.2k steps, to ensure a non cherry-picked results.
Table 3: The pass@$k$ performance of Llama 3.1 8B Instruct on CodeContests and TACO with GRPO / Dr. GRPO objective. Models are trained on CodeContests training set. The pass@$k$ objectives optimize for $k=8$. Pass@$k$ training consistently shows better pass@10 performance across different policy gradient variants.
Variant CodeContests / Valid CodeContests / Test TACO/easy TACO/hard
pass@1 pass@10 pass@1 pass@10 pass@1 pass@10 pass@1 pass@10
mean (Dr. GRPO) 7.17 10.61 8.88 14.89 16.47 22.91 5.40 6.89
mean (GRPO) 9.21 13.88 8.39 14.41 17.89 24.11 4.74 7.06
biased pass@$k$ (Dr. GRPO) 6.79 15.71 7.03 16.44 15.87 27.01 4.86 8.55
biased pass@$k$ (GRPO) 8.11 16.83 6.67 15.45 14.51 26.08 4.12 7.37
Llama 3.1 8B Instruct Performance.
A.3 Training on HARP
@@ B.2 Interpolation between $k$-sample objectives and mean objective
where $S_{p}$ is the set of all subsets of $p$ indices from $\{1,\dots,k\}$, with $\left({k\atop p}\right)$ being its cardinality. Notice that for $p=0$ we recover the initial problem of the pass@$k$ objective. For $p=1$ we get the de-meaned objective introduced in the previous section. For p=k1p=k-1 we get the problem of maximizing the expected rewards. For $p=2$ we get the objective